kenzie126:

Absolute Value Equation: A group of students ran a mile with an average time of 9 minutes. Individual times were no more than 3 minutes higher or lower than the average. Write an absolute value equation to model this situation, and solve the equation to determine the range of times.

jimthompson5910:

Let x be a student's mile time in minutes The distance from x to the average is |x-9| since absolute value is used to measure distance on a number line. Let's say x = 7 minutes, this means |x-9| = |7-9| = 2 tells us the time value x = 7 is two units away from the average. We're told that the highest and lowest times were no further than 3 units away from the average. This must mean we have the absolute value equation |x-9| = 3 Solving this equation for x leads to the highest and lowest times possible.

kenzie126:

i dont know how to solve it

jimthompson5910:

You use the rule that |x| = k turns into x = -k or x = k

jimthompson5910:

In this case, |x-9| = 3 turns into the two equations x-9 = 3 or x - 9 = -3

kenzie126:

woahh what. i dont understand this

jimthompson5910:

what do you get when you solve x-9 = 3

kenzie126:

12?

jimthompson5910:

yes, and how about x-9 = -3 ?

kenzie126:

6?

jimthompson5910:

yes, so 6 is the lowest time and 12 is the highest time We can see this with a number line 

SemiDefinite:

sure?! the Qu posits that: \(\bar t = 9 = \dfrac{\Sigma t_i}{N}\) with: \(\bar t - 3 < t_i < \bar t + 3 \qquad \forall t\)